### 07 Feb 2024 v2

@@ 4 Experiment
We conduct comprehensive experiments across image classification, VQA, and visual grounding tasks within the RS domain. These experiments are designed to rigorously validate the superior multi-task solving capabilities of LHRS-Bot. We present both quantitive and quality results.
@@ 4.1 Experimental Setup
we select diverse RS datasets for providing a comprehensive assessment across different tasks within the RS field: seven RS classification datasets (namely AID [62], WHU-RS19 [11], NWPU [7], SIRI-WHU [79], EuroSAT [17], METER-ML [77], and fMoW [10]), two VQA datasets (LR and HR subsets of RSVQA [35]), and two visual grounding datasets (RSVG [50] and DIOR-RSVG [70]). We exclude any data from the test sets that overlap with the training set to avoid data leakage. More details about each test dataset are presented in Tab. 7.
@@ 4.2 Quantitative Evaluation
We evaluate different MLLMs using greedy decoding with the same prompt for each task and report the average result of two independent trials.
@@
The classification results across different datasets are presented in Tab. 1. These results show that LHRS-Bot outperforms all competing models in all seven classification datasets, with a notable margin of 27.51% higher average accuracy than the second ranked method. It is worth mentioning that, except for NWPU, METER-ML, and fMoW, the remaining datasets are completely absent from the multi-task training dataset  [^4] However, we have no means of confirming whether these data appeared in the LHRS-Align dataset, considering that images in LHRS-Align, along with several of the classification datasets (AID, WHU-RS19, SIRI-WHU), are collected from Google Earth., demonstrating the strong generalization ability of LHRS-Bot. In addition, in the context of the fine-grained RS classification dataset fMoW, our method significantly surpasses other models, leading by an impressive 40% on average. This highlights the critical role of continue pre-training in the RS domain and the effectiveness of our alignment dataset LHRS-Align.
@@ 4.4 Ablation Analysis
4.4.2 Effectiveness of Model Design.
We evaluate the effectiveness of designs involving multi-level visual hidden representations as well as descending query allocation strategies. Specifically, our evaluation focuses on employing the single level visual hidden state, projecting it through a two-layer MLP or a vision perceiver. We also conduct comparisons between multi-level visual hidden representations and incremental query allocation strategies. Building on the concept introduced by the token merging strategy [5], we explore the potential of integrating this strategy with a vision perceiver to efficiently leverage multilevel features. To minimize experimental costs, our training process comprises just two stages: firstly, we employ LHRS-Align to train the bridge layer, which aligns the vision and language domains. Subsequently, we train the model using the LHRS-Instruct and multi-task instruction datasets.
We report the average result on each task in Tab. 5. It is evident that the approach employed by LHRS-Bot achieves superior performance. The results further illustrate that, although the extraction of a singular visual representation performs better in tasks such as classification and VQA tasks, which depend on high-level visual semantics, it falls short in visual grounding tasks that require detailed visual information. Furthermore, employing the same number of queries to summarise different levels of visual information does not lead to better results. In contrast, adopting a diminishing allocation strategy yields superior results. This further demonstrates the large redundancy among the deep visual tokens [5]. Lastly, the token merging strategy proves ineffective, possibly due to the loss of crucial information during the merging process. We leave the investigation and improvement to further works.
@@ 5 Related Work
5.1 Multimodal Large Language Models
With the remarkable capabilities of LLMs in natural language processing [6, 9, 18, 73, 59, 53] and their successful applications across various domains [74, 65, 60, 68, 55, 52, 48, 49, 31, 13], LLMs have been extended to MLLMs by aligning visual inputs with LLMs and further empowered by visual instruction tuning to address a serial of cross-modal tasks [1, 29, 15, 28, 78, 3]. The progression in MLLMs has catalyzed the creation of domain-specific MLLMs, encompassing areas such as biomedicine [28], autonomous driving [37], and remote sensing [25, 71].
5.2 Remote Sensing Vision-Language Models
Early RS vision-language models were designed for specific tasks, such as VQA [35], visual grounding [70], and scene classification [32]. RSGPT [21] is the initial work which exploits LLMs for RS vision-language tasks. However, it necessitates specific fine-tuning for each task, limiting its potential for holistic RS image understanding. Our concurrent works, GeoChat [25] and SkyEyeGPT [71] introduce RS-specific MLLMs which can also address various RS understanding tasks and follow human instructions. However, they encounter challenges to fit existing public RS datasets, due to not fully utilizing extensive and global RS features to enrich LLMs with comprehensive RS visual knowledge. Moreover, they neglect the importance of multi-level visual features for a thorough understanding of RS images, thus restricting their effectiveness in a wide range of tasks at different granularity.
5.3 Large-Scale Remote Sensing Vision-Language Datasets
The RS5M dataset [75], comprised of 5 million RS images, is the first large-scale RS image-text paired dataset. This dataset is curated by filtering publicly available general datasets. Despite the extensive scale, it falls short in providing orthorectified satellite images and offers insufficient geographical descriptions of scenes and objects. SkyScript [57], a dataset developed concurrent with ours, utilizes open-source geographical data to create a large-scale image-text dataset. However, their image captions, created using predefined rules and not thoroughly refined by a rigorous cleaning pipeline, fall short in providing high-quality and rich visual semantics. This limitation hampers the ability to equip LLMs with a wealth of RS visual knowledge.
@@ 6 Conclusion
We propose LHRS-Bot, an MLLM for the RS domain. In our effort to unleash the full potential of LLM for RS image understanding, we curate a specialized large-scale alignment dataset, LHRS-Align, designed for RS-specific alignment, along with LHRS-Instruct, a RS multimodal instruction-following dataset, to facilitate the instruction following ability of LHRS-Bot. In comparative analyses, LHRS-Bot stands out for its exceptional RS domain knowledge, significantly outperforming both the general-purpose and RS-specific MLLMs.
@@
Appendix 0.A Details about LHRS-Align Dataset
0.A.1 Data Source.
For now, we have collected images that cover the 100 most populous cities from certain countries. Due to challenges such as the scarcity of OSM data for certain cities (particularly in underdeveloped regions), the remaining images cover 9,259 cities across 129 countries, as shown in Fig. 6. Considering the spatial resolution limitation of GE images, we only utilize the polygon features of OSM for the current version of our dataset, with the extension of polyline and point features left to our future works. Furthermore, we discard the polygon features tagged with the keys “boundary” and “barrier”, as they do not exactly describe the function of the feature.
@@ Appendix 0.A Details about LHRS-Align Dataset
0.A.2 Geo-alignment and Image Processing.
@@
0.A.5 Details about Generating Captions.
@@
Appendix 0.B Details about LHRS-Instruct Dataset
0.B.1 Instruction-Following Dataset from Public RS Caption Dataset
0.B.1.1 Data Filtering Procedure.
@@ 0.B.1 Instruction-Following Dataset from Public RS Caption Dataset
0.B.1.2 Instruction-Following Data Generation.
We deploy the Vicuna-v1.5-13B [8] for generation with the same setting described in Sec. 0.A.5. The in-context examples are shown in Tab. 11.
0.B.2 Instruction-Following Dataset from LHRS-Align
@@
Appendix 0.C Multi-task Instruction Dataset
@@
Appendix 0.D Details about Test Datasets
@@
Appendix 0.E Hyperparamters

### 18 Mar 2024 v3

@@ 4 Experiment
We conduct comprehensive experiments across image classification, VQA, and visual grounding tasks within the RS domain. These experiments are designed to rigorously validate the superior multi-task solving capabilities of LHRS-Bot. Furthermore, we evaluate the performance of various LLMs in the RS domain using the proposed LHRS-Bench benchmark. We present both quantitive and quality results.
@@ 4.1 Experimental Setup
We select diverse RS datasets for providing a comprehensive assessment across different tasks within the RS field: seven RS classification datasets (namely AID [68], WHU-RS19 [13], NWPU [9], SIRI-WHU [86], EuroSAT [20], METER-ML [84], and fMoW [12]), two VQA datasets (LR and HR subsets of RSVQA [41]), and two visual grounding datasets (RSVG [56] and DIOR-RSVG [77]). We exclude any data from the test sets that overlap with the training set to avoid data leakage. More details about each test dataset are presented in supplementary material.
@@ 4.2 Quantitative Evaluation
We evaluate different MLLMs using greedy decoding with the same prompt for each task and report the average result of two independent trials. Details about the evaluation prompts are provided in supplementary material.
@@
The classification results across different datasets are presented in Tab. 2. These results show that LHRS-Bot outperforms all competing models in all seven classification datasets, with a notable margin of 27.51% higher average accuracy than the second ranked method. It is worth mentioning that, except for NWPU, METER-ML, and fMoW, the remaining datasets are completely absent from the multi-task training dataset  [^6] However, we have no means of confirming whether these data appeared in the LHRS-Align dataset, considering that images in LHRS-Align, along with several of the classification datasets (AID, WHU-RS19, and SIRI-WHU), are collected from Google Earth., demonstrating the strong generalization ability of LHRS-Bot. In addition, in the context of the fine-grained RS classification dataset fMoW, our method significantly surpasses other models, leading by an impressive 40% on average. This highlights the critical role of continue pre-training in the RS domain and the effectiveness of our alignment dataset LHRS-Align.
@@
4.3 Evaluation on LHRS-Bench
We evaluate different MLLMs on our LHRS-Bench benchmark. For each question within the benchmark, choices are shuffled, and the generation process is repeated four times. An answer is considered incorrect if any attempt yields an incorrect response. Same prompts are used for evaluation, as detailed in supplementary material. We consider the model answers correctly if it outputs the correct choice or the answer corresponding to the correct choice. We calculate accuracy as our evaluation metric.
The results, presented in Tab. 5, show that LHRS-Bot achieves the best results on this benchmark. It is worth noting that, LHRS-Bot outperforms others in terms of the evaluation dimensions of spatial resolution and image modality, even though these aspects were not explicitly trained for in the models, highlighting the effectiveness of RS-specific MLLM training. However, although LHRS-Bot has achieved the highest performance, there is still significant room for improvement in MLLM capabilities within the RS domain. Examples of evaluation results are available on supplementary material.
Table 5: Comparison results on the LHRS-Bench dataset.
Method Identity Color Orientation Shape Area Resolution Modality Location Distance Quantity Reasoning Avg.
LLaVA-1.5 42.90 30.97 15.38 43.24 44.00 23.80 21.74 35.78 27.27 21.89 54.35 37.45
MiniGPTv2 20.19 15.93 2.56 8.11 13.33 14.29 4.35 15.20 13.64 9.49 36.96 16.88
InstructBLIP 20.35 22.12 5.13 24.32 22.67 38.10 00.00 13.73 22.85 13.14 30.43 18.87
mPLUG-OWL2 41.48 45.13 12.82 40.54 32.00 33.33 00.00 30.88 13.64 32.12 54.34 37.00
QWen-VL-Chat 31.38 32.74 7.69 32.43 29.33 19.05 4.35 24.02 21.45 21.17 56.52 28.65
InternLM-XComposer 34.38 30.97 7.69 37.84 26.67 00.00 8.70 33.33 22.73 18.25 65.22 31.09
LHRS-Bot 44.64 42.48 23.08 67.57 25.33 42.86 32.35 56.52 22.73 18.25 73.91 39.38
@@ 4.5 Ablation Analysis
4.5.2 Effectiveness of Model Design.
We evaluate the effectiveness of designs involving multi-level visual hidden representations as well as descending query allocation strategies. Specifically, our evaluation focuses on comparing with the single level visual hidden state, projecting it through a two-layer MLP or a vision perceiver. We also perform comparisons between multi-level visual hidden representations, employing either identical queries for each level or incremental query allocation strategies. Building on the concept introduced by the token merging [7], we explore the potential of integrating this strategy with a vision perceiver to efficiently leverage multi-level representations. To minimize experimental costs, our training process comprises just two stages: firstly, we employ LHRS-Align to train the bridge layer, which aligns the vision and language domains. Subsequently, we train the model using the LHRS-Instruct and multi-task instruction datasets.
@@
We report the average result on each task in Tab. 7. It is evident that the approach employed by LHRS-Bot achieves superior performance. The results further illustrate that, although the extraction of a singular visual representation performs better in tasks such as classification and VQA tasks, which depends on high-level visual semantics, it falls short in visual grounding tasks that require detailed visual information. Furthermore, employing the same number of queries to summarise different levels of visual information does not lead to better results. In contrast, adopting a diminishing allocation strategy yields superior results. This further demonstrates the large redundancy among the deep visual tokens [7]. Lastly, the token merging strategy proves ineffective, possibly due to the loss of crucial information during the merging process.
@@ 5 Related Work
MLLMs, specifically vision-language MLLMs, have extended the abilities of LLMs by incorporating visual alignment and visual instruction tuning [1, 34, 17, 33, 85, 3], enabling them to tackle cross-modal tasks in various domains [81, 71, 66, 74, 61, 58, 54, 55, 36, 15, 33, 43]. Concurrently, advancements in RS vision-language models, such as RSGPT [24], GeoChat [28], and SkyEyeGPT [78], have achieved promising improvements in understanding RS images. However, the present RS-specific MLLMs, fine-tuned using limited public task-specific datasets, neglect the potential of leveraging global RS images and vast crowdsourced databases to enhance LLMs with extensive RS visual knowledge. In contrast, we unleash the full potential of LLMs for RS applications by leveraging the abundant geographical information with global available RS images and employing a more effective multi-level alignment strategy. Further details about recent related development are available in supplementary material.
@@ 6 Conclusion
We propose LHRS-Bot, an MLLM for the RS domain. In our effort to unleash the full potential of LLM for RS image understanding, we curate a specialized large-scale alignment dataset, LHRS-Align, designed for RS-specific alignment, along with LHRS-Instruct, a RS multimodal instruction-following dataset, to facilitate the instruction following ability of LHRS-Bot. Furthermore, we establish a benchmark, LHRS-Bench, which provides a systematic evaluation framework for RS MLLMs. In comparative analyses, LHRS-Bot stands out for its exceptional RS domain knowledge, significantly outperforming both the general-purpose and RS-specific MLLMs.
@@
Appendix 0.A Related Work
0.A.1 Multimodal Large Language Models
With the remarkable capabilities of LLMs in natural language processing [8, 11, 21, 80, 65, 59] and their successful applications across various domains [81, 71, 66, 74, 61, 58, 54, 55, 36, 15], LLMs have been extended to MLLMs by aligning visual inputs with LLMs and further empowered by visual instruction tuning to address a serial of cross-modal tasks [1, 34, 17, 33, 85, 3]. The progression in MLLMs has catalyzed the creation of domain-specific MLLMs, encompassing areas such as biomedicine [33], autonomous driving [43], and RS [28, 78].
0.A.2 Remote Sensing Vision-Language Models
Early RS vision-language models were designed for specific tasks, such as VQA [41], visual grounding [77], and scene classification [37]. RSGPT [24] is the initial work which exploits LLMs for RS vision-language tasks. However, it necessitates specific fine-tuning for each task, limiting its potential for holistic RS image understanding. Our concurrent works, GeoChat [28] and SkyEyeGPT [78] introduce RS-specific MLLMs which can also address various RS image understanding tasks and follow human instructions. However, these models are struggled to adapting to existing public RS datasets, thus failing to utilize the extensive and global RS images to enrich LLMs with comprehensive RS visual knowledge. In contrast, we unleash the full potential of LLMs for RS applications by leveraging the abundant geographical information within global available RS images and employing a more effective multi-level alignment strategy.
0.A.3 Large-Scale Remote Sensing Vision-Language Datasets
The RS5M dataset [82], comprised of 5 million RS images, is the first large-scale RS image-text paired dataset. This dataset is curated by filtering publicly available general datasets. Despite the extensive scale, it falls short in providing orthorectified satellite images and offers insufficient geographical descriptions of scenes and objects. SkyScript [63], a dataset developed concurrent with ours, utilizes open-source geographical data to create a large-scale image-text dataset. However, their image captions, created using predefined rules and not thoroughly refined by a rigorous cleaning pipeline, fall short in providing high-quality and rich visual semantics. In this work, we develop a large-scale vision-language RS alignment dataset utilizing the semantically abundant VGI database and global RS images. By implementing a rigorous data curation pipeline and employing LLMs to generate meaningful descriptions from feature tags, the proposed dataset is proven more informative for pretraining an MLLM.
Appendix 0.B Details about LHRS-Align Dataset
0.B.1 Data Source.
For now, we have collected images that cover the 10000 most populous cities from certain countries. Due to challenges such as the scarcity of OSM data for certain cities (particularly in underdeveloped regions), the remaining images cover 9,259 cities across 129 countries, as shown in Fig. 6. Considering the spatial resolution limitation of GE images, we only utilize the polygon features of OSM for the current version of our dataset, with the extension of polyline and point features left to our future works. Furthermore, we discard the polygon features tagged with the keys “boundary” and “barrier”, as they do not exactly describe the function of the feature.
@@ Appendix 0.B Details about LHRS-Align Dataset
0.B.2 Geo-alignment and Image Processing.
@@
0.B.5 Details about Generating Captions.
@@
Appendix 0.C Details about LHRS-Instruct Dataset
0.C.1 Instruction-Following Dataset from Public RS Caption Dataset
0.C.1.1 Data Filtering Procedure.
@@ 0.C.1 Instruction-Following Dataset from Public RS Caption Dataset
0.C.1.2 Instruction-Following Data Generation.
We deploy the Vicuna-v1.5-13B [10] for generation with the same setting described in Sec. 0.B.5. The in-context examples are shown in Tab. 13.
0.C.2 Instruction-Following Dataset from LHRS-Align
@@
Appendix 0.D Multi-task Instruction Dataset
@@
Appendix 0.E Details about Test Datasets
@@
Appendix 0.F Details about LHRS-Bench Dataset
0.F.1 Evaluation Dimensions
The hierarchical ability taxonomies are comprised of 5 top-level dimesntions with 11 fine-grained sub-dimensions. Note that each question-answer pair may encompass multiple sub-dimensions [75]. Detailed introduction of each ability dimensions is outlined below.
  • $\bullet$
    Recognition: This dimension focuses on the model’s ability to recognize instance identity (geographical objects or scene information), and their attributes, including color, orientation, shape, and area.
  • $\bullet$
    Quantity: This dimension assesses the model’s capacity to count specific objects, often combined with a prerequisite understanding of those objects (Recognition).
  • $\bullet$
    Imagery: This dimension centers on the image itself, specifically recognizing its resolution and modality. In the RS domain, images can encompass various modalities, including optical, panchromatic, synthetic aperture radar (SAR), and thermal infrared images. However, as RS MLLMs are still in their early stages of development, LHRS-Bench currently only includes multispectral and panchromatic images.
  • $\bullet$
    Spatial-awareness: This dimension evaluates the MLLM’s ability to comprehend the locations (and spatial relationships between objects), as well as identify the distance between objects.
  • $\bullet$
    Reasoning: This dimension assesses the model’s proficiency in accurately comprehending the image and utilizing its domain-specific knowledge to provide correct answers to questions.
Appendix 0.G Hyperparamters
@@
Appendix 0.H Evaluation Prompt for Each Task
In tasks related to visual grounding and VQA, we prompt each model using descriptions of the target object and questions, respectively. For classification tasks and evaluations on LHRS-Bench, the same prompts are employed across all models, with the specific prompts presented in Tab. 18.
@@ Appendix 0.H Evaluation Prompt for Each Task
Table 18: Prompts for evaluation on each task.
Classification:
Choose the best categories describe the image from: {}.
LHRS-Bench:
Please answer the question based on the given choices:
Question: {}
Choice: {}
Answer:
@@
Figure 9: Examples of the LHRS-Bench benchmark.
Figure 10: Evaluation examples on LHRS-Bench benchmark.